OpenStudy (hba):

Roots of the equation x^4-1=0

OpenStudy (anonymous):

What you can do, is let y = x^2. Then you have y^2 - 1 = 0. Then you can solve that in the usual way, but then realise that you need to convert the y's back into x, through the way you defined y. You'll have 4 solutions, some possibly complex.

OpenStudy (campbell_st):

this of this as \[(x^2)^2 - 1 = 0\] its the difference of 2 squares.

OpenStudy (campbell_st):

so there are 2 real and 2 complex roots